OpenStudy (anonymous):

Solve for x: -3(x - 3) = 21 x = -10 x = 10 x = -4 x = 4

jigglypuff314 (jigglypuff314):

to solve for x: -3(x - 3) = 21 first distribute -3x + 9 = 21 then subtract 9 from both sides divide both sides by negative 3

OpenStudy (anonymous):

4

OpenStudy (jonnyvonny):

Distribute out the -3 toy the both x and -3. -3x+9=21 | -3x=12| x=-4

jigglypuff314 (jigglypuff314):

it was a divide by negative 3 on both sides so you should end up with a negative number :)

OpenStudy (anonymous):

Ok, the first step is to distribute the -3 to (x - 3), this gives you -3x + 9 = 21. Next, you want to isolate the variable. To do this, you need to subtract 9 from 21 giving you -3x = 12. Finally, divide 12 by -3. Giving you a final answer of x = -4